A cylinder with a top and bottom has height 15 in. and has a total surface area that is twice the area of a circle whose radius is 6 in. more than the radius of the cylinder. What is the radius of the cylinder?
Answer by mananth(16946)

A cylinder with a top and bottom has height 15 in. and has a total surface area that is twice the area of a circle whose radius is 6 in. more than the radius of the cylinder. What is the radius of the cylinder?
radius of cylinder=r
radius of circle =r+6
Area of circle= pi * (r+6)^2
2 times area = 2pi(r+6)^2
---
Area of cylinder
2pir^2+2pirh
2pir^2+2pir*15
2pir^2+30pir
equate both
2pir^2+30pir=2pi(r+6)^2
2pir(r+15)=2pi(r+6)^2
r(r+15)=(r+6)^2
r^2+15r=r^2+12r+36
15r-12r=36
3r=36
/3
r=12
radius = 12 inch